In an acid - base titration involving sulphuric acid (H2SO4) and sodium hydroxide (NaOH) solution, 25ml of 0.02m sulphuric acid is required to completely neutralise 20ml of (NaOH) solution.
What is the molarity (mol/L) of NaOH solution?
1
Expert's answer
2018-09-14T05:45:57-0400
Answer:

H2SO4 + 2NaOH = Na2SO4 + 2H2O

Moles of H2SO4 = 25.00*10-3 x 0.02 mol*L-1 = 0.5 x 10-3 mol

[NaOH] = 2*0.5*10-3 mol / 10*10-3 = 0.1 mol*L-1